OpenStudy (anonymous):

How do I identify the slope and y-intercept of the line, ex:8x-2y=8

OpenStudy (espex):

Slope-intercept of a line is \(y=mx+b\), put your equation into that form and you will have your answers.

OpenStudy (anonymous):

so -2y=8x+8

OpenStudy (anonymous):

so that would simplify to y=-4x-4

OpenStudy (anonymous):

so the y intercept would be -4x-4 and the slope would be -4 @eSpeX ?

OpenStudy (espex):

Close, you want to pay closer attention to your signs. \[8x-2y=8\rightarrow-2y=-8x+8\rightarrow y=\frac{-8x}{-2}+\frac{8}{-2}\] and so that would simplify to \(y=4x-4\)

OpenStudy (anonymous):

so y intercept is 4x-4 and slope is 4?

OpenStudy (espex):

In the equation, y=mx+b, 'b' is your y-intercept and m is your slope. So your slope is 4, and your y-intercept is?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

so y intercept is 4x-4 and slope is -4

OpenStudy (espex):

Exactly. Because if 'x' is 0, then you are on the 'y' axis. So if y=4(0)+b, then the y coordinate is y=b

OpenStudy (anonymous):

gotcha thanks!!!!!!!!!!!!!!!

OpenStudy (espex):

You're welcome.

OpenStudy (anonymous):

@eSpeX how would the graph for this look?

OpenStudy (anonymous):

would it be similar to this

OpenStudy (espex):

It would be similar in the sense that the image also shows a line with a positive slope. In your problem you have a slope of 4 and a y-intercept of -4. Since \(m=\frac{rise}{run}\), you have 4=\(\frac{4}{1}\). Starting at your intercept point, (0,-4), you would rise up 4 and run to the right 1 between every point on your line.

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (espex):

You're welcome @aprilsages